OpenStudy (anonymous):

if x1 and x2 are the roots of 3^x-2 + 1= -3^4-x then x1+x2=?

ganeshie8 (ganeshie8):

\( \large 3^{x-2} + 1= -3^{4-x}\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

yes @ganeshie8

ganeshie8 (ganeshie8):

okay isolate the exponent term first

ganeshie8 (ganeshie8):

use below : \(\large a^{m \pm n} = a^m . a^{\pm n} \)

ganeshie8 (ganeshie8):

\(\large 3^{x-2} + 1= -3^{4-x}\) \(\large 3^{x}.3^{-2} + 1= -3^{4}3^{-x}\)

OpenStudy (anonymous):

oky then what should i do?

OpenStudy (anonymous):

will it be \[3^{x+x} = -3^{7-3}\] \[3^{2x} = -3^{4}\]

OpenStudy (anonymous):

oky!!!

ganeshie8 (ganeshie8):

\(\large 3^{x-2} + 1= -3^{4-x}\) multiply thru by 3^2 : \(\large 3^{x} + 3^2 = -3^{6-x}\) multiply again by 3^x : \(\large \left(3^{x}\right)^2 + 3^2\left(3^x\right)= -3^{6}\)

ganeshie8 (ganeshie8):

\(\large \left(3^{x}\right)^2 + 3^2\left(3^x\right)+ 3^{6}=0\)

ganeshie8 (ganeshie8):

Notice that this is a quadratic in 3^x product of roots = c/a = ?

OpenStudy (anonymous):

will it be \[\frac{ 3^{2} }{ (3x)^{2} }\]

ganeshie8 (ganeshie8):

nope, may be let 3^x = t for the time being

ganeshie8 (ganeshie8):

then the equation becomes : \(\large t^2 + 3^2t + 3^6 = 0\)

ganeshie8 (ganeshie8):

product of roots = \(\large t_1 t_2 = 3^6\)

ganeshie8 (ganeshie8):

\(\large \implies \rm 3^{x_1}3^{x_2} = 3^6\)

ganeshie8 (ganeshie8):

\(\large \implies \rm 3^{x_1 + x_2} = 3^6\)

ganeshie8 (ganeshie8):

\(\large \rm \implies x_1 + x_2 = 6\)

ganeshie8 (ganeshie8):

see if that makes more or less sense..

OpenStudy (anonymous):

ohhh yaaa!!!